Direct Labor Variances La Batre Bicycle Company manufactures commuter bicycles from recycled materials. The following data for July of the current year are available: Quantity of direct labor used 640 hrs. Actual rate for direct labor $13.7 per hr. Bicycles completed in April 300 bicycles Standard direct labor per bicycle 2 hrs. Standard rate for direct labor $14 per hr. a. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number. Direct Labor Rate Variance $ -192 Favorable Direct Labor Time Variance $ Unfavorable Total Direct Labor Cost Variance $ Unfavorable b. How much direct labor should be debited to Work in Process? $
Answer #1
a
Direct Labor Rate Variance -192 Favorable =640*(13.7-14)
Direct Labor Time Variance 560 Unfavorable =14*(640-300*2)
Total Direct Labor Cost Variance 368 Unfavorable =(640*13.7)-(300*2*14)
b
Direct labor debited to Work in Process 8400 =300*2*14